# Recovery of Sparse Graph Signals

### v1

@@ III Dictionary Matrix Analysis
Due to the graph filter structure defined by (6)-(8), and the sparse connectivity of the graph imposed by Assumption A.1, we can expect a decrease in the energy of the projection of ${{\bf{H}}}_{{\Omega}}{{\bf{x}}}_{{\Omega}_{t}}$ over the elements in ${\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}$ as the distance from the elements in ${\Omega}_{t}$ increases. Thus, there exists a set $\mathcal{D}\subset\mathcal{N}_{2\Psi}({\Omega}_{t})$ such that
$\|{\bf{P}}_{{{\bf{H}}}_{m}}{{\bf{H}}}_{{\Omega}_{t}}{{\bf{x}}}_{{\Omega}_{t}}\|_{2}^{2}<\epsilon,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall m\in\{\mathcal{N}_{2\Psi}({\Omega}_{t})\setminus\mathcal{D}\},$ (29)
where $\epsilon$ is small. Therefore, it is implied that the dictionary matrix elements that contain information on the filtered signal is included in the set $\mathcal{D}$, which is a subset of the $2{\Psi}$-hop neighborhood of ${\Omega}_{t}$, $\mathcal{N}_{2{\Psi}}[{\Omega}_{t}]$, $\mathcal{D}\subset\mathcal{N}_{2{\Psi}}[{\Omega}_{t}]$. As a result, when the sparse signal input is distributed across different regions of the graph, the nodes within ${\mathcal{D}}$, which are composed of only a small portion of the graph nodes, will also tend to be in distant regions of the graph.
@@ IV Sparse Recovery of Graph Signals
In this section, we propose two methods for solving the GIC support set recovery problem in (18) under the GSP measurement model in (11). The analysis in Section III provides the basis for these methods, since it enables searching for the optimal support set over a reduced set of dictionary atoms. In Subsection IV-A, we present the GM-GIC method, which involves a node partitioning followed by the implementation of the GIC locally on each partitioned subset. In Subsection IV-B, we present the Graph-BNB-GIC method, which is based on the BNB method [60, 61], and involves an upper bound that is derived based on the underlying graphical structure, thus significantly enhancing the traceability of this method.
IV-A GM-GIC
In this subsection, we outline the rationale and then present the methodology of the GM-GIC method.
IV-A1 Rationale
The analysis in Section III implies that the true support set of the signal, ${\Omega}_{t}$, is included in the set ${\mathcal{D}}$ from (29). Considering that ${\lvert{\mathcal{D}}\rvert}\ll{\lvert{\mathcal{V}}\rvert}$, in many cases the elements in ${\mathcal{D}}$ are located in different regions across the graph as discussed below (29). Thus, in many cases the set $\mathcal{D}$ can be divided into $Q$ subsets, $\{\mathcal{D}_{q}\}_{q=1}^{Q}$, such that
@@ 2 Set                  % new edge set
3 Generate a partition of $\tilde{\mathcal{G}}(\tilde{{\mathcal{V}}},\tilde{{\mathcal{E}}})$ with connected components (e.g. using conncomp in Matlab)
@@ IV-A2 Methodology
The GM-GIC method consists of four steps. In the first step, based on (23), we estimate the set $\mathcal{D}$ described in (29) by applying the generalized likelihood ratio test (GLRT) for each $m\in\mathcal{V}$ on the following binary hypothesis testing problem:
@@
where $\zeta$ is a user-defined threshold parameter. Thus, the set ${\mathcal{D}}$ is estimated according to the binary detector in (35), which is defined for each $m\in{\mathcal{V}}$, by
$\hat{{\mathcal{D}}}\stackrel{{\scriptstyle\triangle}}{{=}}\{m\in\mathcal{V}:\leavevmode\nobreak\ \|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{y}}}\|^{2}>\zeta\}.$ (36)
@@
$\hat{{\Omega}}_{q}=\arg\underset{\{\lvert{\Omega}\rvert\leq{s},{\Omega}\subseteq\hat{\mathcal{D}}_{q}\}}{\max}\leavevmode\nobreak\ \|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}}}}}{{\bf{y}}}\|^{2}-{\rho({\lvert{\Omega}\rvert})},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ q=1,\ldots,Q.$ (37)
@@
$\hat{{\Omega}}^{temp}=\bigcup\limits_{q=1}^{Q}\hat{{\Omega}}_{q}.$ (38)
@@
$\hat{{\Omega}}=\arg\underset{\{\lvert{\Omega}\rvert\leq{s},{\Omega}\subseteq\hat{{\Omega}}^{temp}\}}{\max}\leavevmode\nobreak\ \|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}}}}}{{\bf{y}}}\|^{2}-{\rho({\lvert{\Omega}\rvert})}.$ (39)
The GM-GIC method is summarized in Algorithm 2 and illustrated in Fig. 1.
@@ 4       Obtain $\hat{{\Omega}}_{q}$ by solving (37)
Figure 1: Steps 1-3 of the GM-GIC method from Algorithm 2. In Step $1$, the algorithm estimates the candidate node set, ${\mathcal{D}}$, denoted by black-filled nodes. A node $k$ is included in ${\mathcal{D}}$ if the projection of the measurements into the signl-column space $\text{col}({{\bf{H}}}_{k})$, $\lvert\lvert{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{k}}}}{\bf{y}}\rvert\rvert^{2}$, exceeds a certain threshold. In Step $2$, the nodes in ${\mathcal{D}}$ are partitioned into subsets (here, 2 subsets in pink and green) based on the geodesic distance between the nodes, the graph filter order, and the graph structure. Finally, in Step $3$, a local GIC is performed for the support set recovery within the subsets. This results in separate partial support sets (denoted by red rings around the nodes).
IV-B Graph-BNB-GIC
In this subsection we present a new Graph-BNB-GIC method for solving the GIC problem in (18). To this end, we reformulate the GIC problem in (18) as an integer programming optimization problem, and then derive the Graph-BNB-GIC method. The Graph-BNB-GIC, is a BNB-based method (see, e.g. [60, 61]) with novel graph-based designs for the branching and bounding steps.
@@ IV-B1 New formulation
By using the ${\lvert{\mathcal{V}}\rvert}\times 1$ Boolean vector ${{\bf{z}}}=[{z_{1}},\ldots,{z_{{\lvert{\mathcal{V}}\rvert}}}]^{T}$, we adress the GIC problem in (18) as the following integer programming optimization problem:
$\hat{{{\bf{z}}}}=\arg\underset{{{\bf{z}}}\in\{0,1\}^{{\lvert{\mathcal{V}}\rvert}}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}({{\bf{z}}})}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert{\Omega}({{\bf{z}}})\rvert})}$
@@ IV-B2 Branch
According to the discussion around (29), nodes included in the true support set of the signal ${{\bf{x}}}$, ${\Omega}_{t}$, are associated with smaller indices in (41). Moreover, dictionary atoms associated with nodes indexed larger than ${\lvert{\mathcal{D}}\rvert}$, w.r.t. to the ordering in (41), are not expected to contain substantial information on the filtered signal, ${{\bf{H}}}{{\bf{x}}}$. Based on the ordering in (41), we describe the branch step as follows. By selecting index $1$, we form the following two problems: the first problem is
$\hat{{{\bf{z}}}}=\arg\underset{{{\bf{z}}}\in\{0,1\}^{{\lvert{\mathcal{V}}\rvert}}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}({{\bf{z}}})}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert{\Omega}({{\bf{z}}})\rvert})}$
@@
$\hat{{{\bf{z}}}}=\arg\underset{{{\bf{z}}}\in\{0,1\}^{{\lvert{\mathcal{V}}\rvert}}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}({{\bf{z}}})}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert{\Omega}({{\bf{z}}})\rvert})}$ (43)
@@
$\hat{{{\bf{z}}}}=\arg\underset{{{\bf{z}}}\in\{0,1\}^{{\lvert{\mathcal{V}}\rvert}}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}({{\bf{z}}})}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert{\Omega}({{\bf{z}}})\rvert})}$
@@ IV-B3 Bound
${\boldsymbol{\Phi}}_{ub}(\mathcal{S}_{0},\mathcal{S}_{1})={\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{\mathcal{S}_{1}}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert\mathcal{S}_{1}\rvert})}\hskip 71.13188pt$
@@
    The cost function associated with the elements in ${\bf{z}}$ that are already fixed as $1$ (the elements in $\mathcal{S}_{1}$): ${\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{\mathcal{S}_{1}}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert\mathcal{S}_{1}\rvert})}$;
@@
${\boldsymbol{\Phi}}_{lb}(\mathcal{S}_{1})={\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{\mathcal{S}_{1}}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert\mathcal{S}_{1}\rvert})}.$ (47)
@@ IV-B4 Remarks
$+(s-{\lvert\mathcal{S}_{1}\rvert})\max\left({\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{k+1}}}}{{\bf{y}}}\|}^{2}-\rho(1),0\right).$ (48)
@@
Due to the ordering in (41), it can be verified that if $k>{\lvert{\mathcal{D}}\rvert}$, then $k\in\{\mathcal{V}\setminus{\mathcal{D}}\}$. Hence, from (29) we obtain that ${\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{k}}}}{\bf{H}}_{{\Omega}_{t}}{{\bf{x}}}_{{\Omega}_{t}}\|}<\epsilon$. Thus, by considering a low degree of noise, it can be assumed that
${\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{k+1}}}}{{\bf{y}}}\|}^{2}-\rho(1)<\epsilon.$ (49)
By substituting the result from (49) in (1), we obtain that the difference between the upper and lower bound in (1) when $k>{\lvert{\mathcal{D}}\rvert}$ satisfies
@@
where the sparsity parameter ${s}$ indicates the the sparsity level of the signal, ${{\bf{x}}}$, and $\epsilon$ is defined in (29). This result implies that observing support set elements with nodes outside ${\mathcal{D}}$ cannot improve the support set recovery results. Consequently, the result shows that in the worst-case scenario of the algorithm, it terminates after observing all support set combinations within the set $\mathcal{D}$ from (29). Hence, in its worst case, the algorithm requires an order of $O\left(\sum_{j=1}^{{s}}\binom{|\mathcal{D}|}{j}\right)$ flops (see Subsection IV-C for a comparison with the GIC and GM-GIC methods).

### v2

@@ III Dictionary Matrix Analysis
Dictionary atoms within the set $\mathcal{N}_{2\Psi}({\Omega}_{t})$ may differ in the significance of the information they provide about the signal. Specifically, some dictionary atoms might convey less relevant information compared to others, with some containing only minor details that are less significant in the context of the signal. To remove the less informative dictionary atoms, we define the subset $\mathcal{D}$ as including only those dictionary atoms that offer substantial information about the signal.
Definition 2.
Let $\epsilon>0$ be a tuning parameter. Then, the set $\mathcal{D}\subset\mathcal{N}_{2\Psi}({\Omega}_{t})$ is defined such that
$\|{\bf{P}}_{{{\bf{H}}}_{m}}{{\bf{H}}}_{{\Omega}_{t}}{{\bf{x}}}_{{\Omega}_{t}}\|_{2}^{2}<\epsilon,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall m\in\{\mathcal{N}_{2\Psi}({\Omega}_{t})\setminus\mathcal{D}\}.$ (31)
The definition of the set $\mathcal{D}$ is determined by the parameter $\epsilon$, introducing a fundamental trade-off. A larger value of $\epsilon$ will result in a smaller set $\mathcal{D}$, which may improve the computational efficiency of the proposed sparse recovery algorithms. However, this can also lead to the exclusion of dictionary atoms containing significant signal information from the set $\mathcal{D}$, thereby degrading the overall recovery performance. Conversely, selecting a smaller value of $\epsilon$ results in a larger set $\mathcal{D}$, potentially reaching the extreme case where $\mathcal{D}=\mathcal{N}_{2\Psi}({\Omega}_{t})$. In this scenario, the size of the search space and the computational overhead are not effectively reduced. In addition, the choice of $\epsilon$ can ensure the existence of a non-empty set $\mathcal{D}$. Thus, careful tuning of $\epsilon$ is crucial. One approach is to set $\epsilon$ based on the snr (snr), ensuring that atoms with contributions comparable to the noise variance are excluded from $\mathcal{D}$. Another strategy is to leverage prior information, if available, to to guide the selection of $\epsilon$.
Using the set $\mathcal{D}$ instead of the entire node set ${\mathcal{V}}$ for the objective function in (18), i.e. searching over $\{{\Omega}\subset\mathcal{D}:\leavevmode\nobreak\ {\lvert{\Omega}\rvert}\leq{s}\}$, offers significant advantages in some scenarios, depending on the sparsity pattern of the signal. In these scenarios, reducing the search space may enable the partitioning of $\mathcal{D}$ into subsets, $\{\mathcal{D}_{q}\}_{q}$ that are associated with dictionary atoms spanning orthogonal subspaces, as detailed in Section IV-A1. In this case, the search set can be divided wrt the partition to smaller search sets: $\left\{\{{\Omega}\subset\mathcal{D}_{q}:\leavevmode\nobreak\ {\lvert{\Omega}\rvert}\leq{s}\}\right\}_{q}$. This strategic partitioning facilitates parallel processing, which enhances the computational efficiency during the search over $\mathcal{D}$.
@@ IV Sparse Recovery of Graph Signals
In this section, we propose two methods for solving the gic support set recovery problem in (18) under the gsp measurement model in (11). The analysis in Section III provides the basis for these methods, since it enables searching for the optimal support set over a reduced set of dictionary atoms. In Subsection IV-A, we present the gmgic method, which involves a node partitioning followed by the implementation of the gic locally on each partitioned subset. In Subsection IV-B, we present the gbnb method, which is based on the bnb method [65, 66], and involves an upper bound that is derived based on the underlying graphical structure, thus significantly enhancing the traceability of this method.
IV-A gmgic
In this subsection, we outline the rationale and then present the methodology of the gmgic method.
IV-A1 Node partition
The analysis in Section III implies that the true support set of the signal, ${\Omega}_{t}$, is included in the set ${\mathcal{D}}$ from (31). We define the following partition on the set $\mathcal{D}$, $\{\mathcal{D}_{q}\}_{q=1}^{Q}$, such that
@@ 2 Set                  % new edge set
Generate a partition of $\tilde{\mathcal{G}}(\tilde{{\mathcal{V}}},\tilde{{\mathcal{E}}})$ with connected components (e.g. using conncomp in Matlab)
@@
The node partitioning presented in this subsection is used in the gmgic method in Subsection IV-A2 in order to reduce the amount of required computations without compromising the estimation performance. When partitioning is not possible (i.e. when $Q=1$), gmgic method still offers an advantage over the gic method by performing an exhaustive search over $\mathcal{D}$, which is a smaller subset of the overall node set $\mathcal{V}$. Moreover, when $Q>1$, partitioning the subset $Q$ into smaller subsets $\{\mathcal{D}_{q}\}_{q}$ enables the exhaustive search to be conducted on these smaller subsets, as illustrated in Fig. 1.
@@ IV-A2 Methodology
The gmgic method consists of four steps. In the first step, based on (24b), we estimate the set $\mathcal{D}$ described in (31) by applying the glrt (glrt) for each $m\in\mathcal{V}$ on the following binary hypothesis testing problem:
@@
where $\zeta$ is a user-defined threshold parameter. This threshold is closely related to the parameter $\epsilon$ from the definition of the set $\mathcal{D}$ in (31). However, while $\epsilon$ is associated with the filtered signal ${{\bf{H}}}{{\bf{x}}}$ (see in (31)), $\zeta$ is applied to the measurement vector ${\bf{y}}$ (see in (37)) and, thus, also accounts for the measurement noise. Thus, the set ${\mathcal{D}}$ is estimated according to the binary detector in (37), which is defined for each $m\in{\mathcal{V}}$, by
$\hat{{\mathcal{D}}}\mathrel{{}\mathop{=}\limits^{\triangle}}\{m\in\mathcal{V}:\leavevmode\nobreak\ \|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{y}}}\|^{2}>\zeta\}.$ (38)
@@
$\hat{{\Omega}}_{q}=\arg\underset{\{\lvert{\Omega}\rvert\leq{s},{\Omega}\subseteq\hat{\mathcal{D}}_{q}\}}{\max}\leavevmode\nobreak\ \|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}}}}}{{\bf{y}}}\|^{2}-{\rho{\lvert{\Omega}\rvert}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ q=1,\ldots,Q.$ (39)
@@
$\hat{{\Omega}}^{temp}=\bigcup\nolimits_{q=1}^{Q}\hat{{\Omega}}_{q}.$ (40)
@@
$\hat{{\Omega}}=\arg\underset{\{\lvert{\Omega}\rvert\leq{s},{\Omega}\subseteq\hat{{\Omega}}^{temp}\}}{\max}\leavevmode\nobreak\ \|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}}}}}{{\bf{y}}}\|^{2}-{\rho{\lvert{\Omega}\rvert}}.$ (41)
The gmgic method is summarized in Algorithm 2 and illustrated in Fig. 1.
@@ 4       Obtain $\hat{{\Omega}}_{q}$ by solving (39)
Figure 1: Steps 1-3 of the gmgic method from Algorithm 2. In Step $1$, the algorithm estimates the candidate node set, ${\mathcal{D}}$, denoted by black-filled nodes. A node $k$ is included in ${\mathcal{D}}$ if the projection of the measurements into the single-column space $\text{col}({{\bf{H}}}_{k})$, $\lvert\lvert{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{k}}}}{\bf{y}}\rvert\rvert^{2}$, exceeds a certain threshold. In Step $2$, the nodes in ${\mathcal{D}}$ are partitioned into subsets (here, 2 subsets in pink and green) based on the geodesic distance between the nodes, the graph filter order, and the graph structure. Finally, in Step $3$, a local gic is performed for the support set recovery within the subsets. This results in separate partial support sets (denoted by red rings around the nodes).
IV-B gbnb
In this subsection we present a new gbnb method for solving the gic problem in (18). To this end, we reformulate the gic problem in (18) as an integer programming optimization problem, and then derive the gbnb method. The gbnb, is a bnb-based method (see, e.g. [65, 66]) with novel graph-based designs for the branching and bounding steps.
@@ IV-B1 New formulation
By using the ${\lvert{\mathcal{V}}\rvert}\times 1$ Boolean vector ${{\bf{z}}}=[{z_{1}},\ldots,{z_{{\lvert{\mathcal{V}}\rvert}}}]^{T}$, we address the gic problem in (18) as the following integer programming optimization problem:
$\hat{{{\bf{z}}}}=\arg\underset{{{\bf{z}}}\in\{0,1\}^{{\lvert{\mathcal{V}}\rvert}}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}({{\bf{z}}})}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert{\Omega}({{\bf{z}}})\rvert}}$
@@ IV-B2 Branch
According to the discussion around (31), nodes included in the true support set of the signal ${{\bf{x}}}$, ${\Omega}_{t}$, are associated with smaller indices in (43). Moreover, dictionary atoms associated with nodes indexed larger than ${\lvert{\mathcal{D}}\rvert}$, wrt to the ordering in (43), are not expected to contain substantial information on the filtered signal, ${{\bf{H}}}{{\bf{x}}}$. Based on the ordering in (43), we describe the branch step as follows. By selecting index $1$, we form the following two problems: the first problem is
$\hat{{{\bf{z}}}}=\arg\underset{{{\bf{z}}}\in\{0,1\}^{{\lvert{\mathcal{V}}\rvert}}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}({{\bf{z}}})}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert{\Omega}({{\bf{z}}})\rvert}}$
@@
$\hat{{{\bf{z}}}}=\arg\underset{{{\bf{z}}}\in\{0,1\}^{{\lvert{\mathcal{V}}\rvert}}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}({{\bf{z}}})}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert{\Omega}({{\bf{z}}})\rvert}}$ (45)
@@
$\hat{{{\bf{z}}}}=\arg\underset{{{\bf{z}}}\in\{0,1\}^{{\lvert{\mathcal{V}}\rvert}}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}({{\bf{z}}})}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert{\Omega}({{\bf{z}}})\rvert}}$
@@ IV-B3 Bound
${\boldsymbol{\Phi}}_{ub}(\mathcal{S}_{0},\mathcal{S}_{1})={\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{\mathcal{S}_{1}}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert\mathcal{S}_{1}\rvert}}\hskip 71.13188pt$
@@
    The cost function associated with the elements in ${\bf{z}}$ that are already fixed as $1$ (the elements in $\mathcal{S}_{1}$): ${\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{\mathcal{S}_{1}}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert\mathcal{S}_{1}\rvert}}$;
@@
${\boldsymbol{\Phi}}_{lb}(\mathcal{S}_{1})={\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{\mathcal{S}_{1}}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert\mathcal{S}_{1}\rvert}}.$ (49)
@@ IV-B4 Remarks
$+(s-{\lvert\mathcal{S}_{1}\rvert})\max\left({\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{k+1}}}}{{\bf{y}}}\|}^{2}-\rho,0\right).$ (50)
@@
Due to the ordering in (43), it can be verified that if $k>{\lvert{\mathcal{D}}\rvert}$, then $k\in\{\mathcal{V}\setminus{\mathcal{D}}\}$. Hence, from (31) we obtain that ${\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{k}}}}{\bf{H}}_{{\Omega}_{t}}{{\bf{x}}}_{{\Omega}_{t}}\|}<\epsilon$. Thus, if the noise variance is negligible compared to the filtered signal, it can be assumed that
${\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{k+1}}}}{{\bf{y}}}\|}^{2}-\rho<\epsilon,$ (51)
where $\rho$ is the gic penalty parameter in (18). By substituting the result from (51) in (1), we obtain that the difference between the upper and lower bound in (1) when $k>{\lvert{\mathcal{D}}\rvert}$ satisfies
@@
where the sparsity parameter ${s}$ indicates the sparsity level of the signal, ${{\bf{x}}}$, and $\epsilon$ is defined in (31). This result implies that observing support set elements with nodes outside ${\mathcal{D}}$ cannot improve the support set recovery results. Consequently, the result shows that in the worst-case scenario of the algorithm, it terminates after observing all support set combinations within the set $\mathcal{D}$ from (31). Hence, in its worst case, the algorithm requires an order of $O\left(\sum_{j=1}^{{s}}\binom{|\mathcal{D}|}{j}\right)$ flops (see Subsection IV-C for a comparison with the gic and gmgic methods).